OpenStudy (anonymous):

A satellite GeoSAT is in a circular geostationary orbit of radius RG, above a point P on the equator. Another satellite ComSAT is in a lower circular orbit of radius 0.81RG. At 7 P.M on January 1, ComSAT is sighted directly above P. On which DATE can ComSAT be sighted directly above P between 7 P.M. and 8 P.M.?

OpenStudy (jamesj):

By one of Kepler's laws, the period T of an orbit and the semi-major axis of that orbit R (or radius if the motion is circular) are proportional according to the following relationship: \[ T^2 \ \alpha \ R^3 \] Call G the geostationary satellite and C the other. Then \[ \frac{T_G^2}{R_G^3} = \frac{T_C^2}{R_C^3} \] Now you know that \[ T_G = 24 \ hours \] precisely because G is geostationary. From this, the other data in the question, and the relationship above, you can deduce the period of C, \( T_C \). And then you can answer the question.

OpenStudy (anonymous):

yahhh thanks.....